# with Knowledge Distillation for Question Answering

### v4

@@ 5.1 Comparison with Human-Annotated Datasets
Unlike prior methods like MultiModalQA (Talmor et al., 2021) and WebQA (Chang et al., 2021), our approach is fully automated with no human involvement (minimal human evaluation was used in the M2QA-Bench only). This section compares the quality of our synthesized data against these human-annotated datasets. We trained LLaVA-1.6 (Liu et al., 2023b; a; 2024), InternVL-2 (Chen et al., 2023b; 2024c), and Idefics-2 (Laurençon et al., 2023; 2024b) on varying sizes of WebQA and MultiModalQA, evaluating on their respective test sets. We also trained the same models on FM2DS-generated training data and evaluated them on the same test sets to assess the effectiveness of the synthesized data in comparison to human-annotated datasets.
As shown in Table 3, models trained on FM2DS data outperform those trained on original datasets, despite using longer documents. WebQA seems easier than MultiModalQA, with better performance from fewer samples. On average, EM improved by 1.81 for MultiModalQA and 1.96 for WebQA using equal or fewer synthetic samples. While EM gains often lead to higher F1, some cases show F1 drops due to hallucinated answers reducing string overlap. Models trained on fewer synthetic samples from FM2DS match the performance of those trained on full datasets, showing faster convergence (see Section 5.2). Larger models also perform better with the same synthetic data—e.g., LLaVA-1.6-13B vs. 7B. GPT-4o leads among large LVLMs, likely due to its role in data generation (Refer to Appendix J for qualitative analysis). When comparing these human annotated data with our synthesized data, a common question is: ”Why not paraphrase existing datasets instead of synthesizing new ones?”. The answer is paraphrasing doesn’t enable domain adaptation, and as shown in Appendix K, fine-tuning on our synthetic data outperforms training on a paraphrased version of MultiModalQA.
@@ 5.2 Learning Efficiency Comparison
To evaluate learning efficiency, we ran experiments with InternVL-2-8B using incremental training sizes from 1k to 10k (in 1k steps) on both synthetic and human-annotated data. For real data, we used the same dataset for training and testing—e.g., when testing on WebQA, training samples were taken from the WebQA training set.
As shown in Figure 3(a & b), our synthesized data outperforms real data at smaller training sizes, though the gap narrows as sample size grows. Near 10k samples, learning efficiency with synthetic data declines more than with real data—likely due to its broader knowledge coverage. While this diversity aids early learning, it can lead to saturation, unlike real data, which offers more focused patterns and sustains steady learning (Hong et al., 2023; Maini et al., 2024).
In a related experiment on MultiModalQA, we used full Wikipedia pages via linked articles as training data instead of information snippets. This wasn’t possible for WebQA, as its source links mostly point to WikiMedia pages with limited text. Figure 3(c) shows that models trained on full Wikipedia pages initially achieve better improvement per 1k samples; however, this advantage declines after approximately 3k samples. This suggests that, while full-page real data offers some early benefits over short snippets, it still lacks the generality and consistency of our synthesized dataset. In contrast, the synthesized data—with its built-in queries and diverse content—continues to support steady learning, likely due to its broader coverage and higher quality.
@@
5.3 Cross-Dataset Evaluation
To assess the generalizability of our synthesized data, we conducted a cross-dataset evaluation using InternVL-2-8B. We used 1k samples from M2QA-Bench and 1k from the MultiModalQA test set, both with full Wikipedia pages as sources. The model was trained separately on 5k samples from our dataset and 5k from MultiModalQA, using the same input format. As shown in Table 4, the model trained on our data outperformed the one trained on MultiModalQA across both test sets, demonstrating stronger generalization. This also reflects the greater complexity and diversity of our benchmark, compared to MultiModalQA’s template-based questions. See Appendix H to see the performance on model trained on FM2DS’s data when using on a out of domain MMQA dataset.
@@ 5.4 Key Stages in FM2DS
Table 5: Results of the FM2DS with and without key steps like query generation and verification. All other steps are included in all of the results. The plus sign (”+”) at the start each steps means all the previous steps were included as well.
To assess the impact of each data filtering and validation step, we evaluated InternVL-2-8B on 5k training samples under different configurations. As shown in Table 5, each step contributed to performance gains. Question validation improved relevance by filtering questions tailored to the complex, multimodal, multihop requirements of the test sets, and reduced hallucinations, boosting F1 with more accurate answers. Answer validation removed incorrect samples, while query generation distilled knowledge from larger models, improving EM and F1. Query validation reinforced consistency by ensuring proper structure. Refer to Appendix L for statistics like ”Rejection Rate” on the validation stages used to remove specific samples during data synthesis.
@@ 5.5 M2QA-Bench Evaluation Results
To assess the complexity of M2QA-Bench and compare model performance, we evaluated a diverse set of models, as shown in Figure 4. GPT-4o stands out, outperforming even larger models like LLaMa-3.2-90B and Claude-3.5-Sonnet by a notable margin. Interestingly, smaller models in the 4B–8B range, particularly those from the Phi family (Abdin et al., 2024a; b), also achieve competitive results despite their scale. GPT-4o’s advantage may stem in part from its involvement in the initial data generation, but this remains an open question, as human annotators thoroughly reviewed and corrected the dataset to ensure fairness, reduce bias, and maintain high quality.
@@ 6 Human Evaluation of Answer Validation
To evaluate the accuracy and impact of our automatic answer validation component, we conducted a human evaluation study. After generating questions for 100 samples, we continued the pipeline using two methods: FM2DS with and without answer validation. These samples were divided into four batches of 25, with each batch evaluated by three participants to mitigate potential bias or human errors.
In total, 12 participants contributed through our evaluation platform (see Appendix F), assessing which method produced the correct answer.
@@
Participants were instructed to verify the correctness of each answer by reviewing the associated Wikipedia pages. Once they determined the accuracy of each answer, they were asked to select one of four options: (1) Method 1’s answer is correct, (2) Method 2’s answer is correct, (3) both methods generated the correct answer, or (4) neither answer is correct (Method 1 and Method 2 were randomly assigned to the datasets generated with and without answer validation, respectively.). The results, shown in Figure 5, reveal a clear trend: answer validation increases the likelihood of correct answers, even though the model without validation still produced correct answers in over 60% of cases. The Fleiss’ Kappa was 0.91—indicating strong agreement—though this is expected, as the task involved factual questions with definitive answers rather than subjective judgments.
@@ 7 Conclusion & Future Works
We recruited human evaluators to support our research and compensated them at a rate of \$10 per hour. All evaluators provided informed consent prior to participation and were informed about the nature of the task. The annotation process adhered to ethical research guidelines, ensuring fair compensation and voluntary participation. No personally identifiable information was collected.
We also recruited three human annotators to verify and clean our M2QA-Bench data. These annotators were compensated at a rate of \$25 per hour for their expertise and time. All annotators provided informed consent and were briefed on the goals and procedures of the task. The annotation process followed ethical research guidelines, ensuring fair compensation, data confidentiality, and voluntary participation throughout.
@@ Appendix B Experimental Settings
Appendix C The Effect of the Number of In-Context Examples
@@
Table 6 presents the results of evaluating the Intervl-2-8B model with varying numbers of in-context examples on the MultiModalQA and WebQA datasets. In the zero-shot setting, FM2DS exhibits limited understanding of multimodal multi-hop question answering, and occasionally circumvents the validation step by simply generating a question that is not multihop. For example, ”Looking at the image of the Eiffel Tower, what engineering innovation allowed it to surpass previous structures in height?” prompts the model to use the image, but the answer is available in the page’s text on tall structures. As we move from zero-shot to one-shot, there is a significant boost in EM and F1 scores, reflecting improved performance with minimal context. The improvement from one-shot to two-shot is marginal, suggesting diminishing returns. With three in-context samples, the gains become minimal, indicating that additional samples beyond two provide little benefit. This diminishing return may stem from the model’s limited context window, which restricts its ability to fully utilize large in-context samples (Kaplan et al., 2020; Bertsch et al., 2024).
Appendix D Comparing Methods For Data Synthesis
@@
Appendix E Investigating FM2DS on Other Models
@@
Appendix F Human Evaluation Details
@@
In addition to these selections, participants had the option to provide a brief rationale for their choices. Although they have not been investigated for this research, this optional feedbacks were encouraged, as it offers valuable insights for qualitative analysis and potential future improvements in answer validation accuracy. The combination of structured and open-ended responses enhances the robustness of our evaluation and offers a more comprehensive view of user judgments, which we may explore in future iterations of our data synthesis methodology.
@@
Appendix G Additional Statistics & Information on M2QA-Bench
To further examine the diversity of questions in our benchmark—which also reflects the overall characteristics of the data generated by FM2DS—we conducted a 2D t-SNE analysis of question embeddings using ModernBERT (Warner et al., 2024). We sampled 500 questions each from M2QA-Bench, MMQA, WebQA, and ScienceQA (Lu et al., 2022a). ScienceQA serves as a fully human-authored dataset, while MMQA and WebQA primarily use templated questions. As shown in Figure 9, MMQA and WebQA display the least diversity. In contrast, M2QA-Bench, which includes questions generated from FM2DS, demonstrates greater similarity to human-generated data, reflecting a reduced domain gap and improved diversity compared to MMQA and WebQA.
@@ Appendix G Additional Statistics & Information on M2QA-Bench
Appendix H FM2DS Generalizability to Other Domains
FM2DS can generate domain-specific synthetic data using just a single in-context example and user-provided documents, enabling even small LVLMs to handle domain-specific multimodal multihop QA. As shown in Figure 7, FM2DS’s data (including M2QA-Bench) spans various domains. To further assess its generalizability, we trained InternVL-2-8B on 5k synthesized samples and evaluated it on the health-related dataset PMC-VQA (Zhang et al., 2023) using 50k test samples. We then compared its performance to the pretrained InternVL-2-8B on the same test set. Training improved accuracy from 52.82 to 65.72, demonstrating the benefits of our synthetic data.
Appendix I M2QA-Bench Examples
Figure 10: Multimodal and multihop reasoning example from M2QA-Bench where the model answers a question about the photograph ”Raising the Flag on Iwo Jima” by synthesizing information from linked documents through a hyperlink, leveraging both visual and tabular data to determine the number of casualties from the USA.
Figure 11: Multimodal multihop reasoning example from M2QA-Bench where the model compares the release dates of two albums, ”Music from Big Pink” and ”Imagine,” using textual and visual cues. The documents are connected through their shared topic, ”music,” and the answer is determined as the title of the earlier-released album.
@@ Appendix I M2QA-Bench Examples
Appendix J Qualitative Analysis
@@
This analysis was conducted for 100 samples from each of the following benchmarks: (1) M2QA-Bench, (2) MultiModalQA, and (3) WebQA. The results are presented in Tables 9, 10, and 11. The responses generated by different models were analyzed across these datasets, focusing on the following metrics:
@@
Table 9: Performance comparison of different model families fine-tuned on real and synthesized data on M2QA-Bench. The ratios for EM scores and hallucination were calculated from filtered data (e.g., hallucination as the proportion of hallucinated responses to incorrect answers). In the table, $\uparrow$ indicates that larger values are better (all EM values), while $\downarrow$ indicates that smaller values are better (hallucination rate). The scores for EM (Table) and EM (Image) may include samples that also contain other modalities. Larger models and those fine-tuned on synthesized data generally show improved performance with reduced hallucination rates. In M2QA-Bench, models demonstrate higher performance on questions involving table modality compared to those involving images.
Model Trained On EM $\uparrow$ Hallucination $\downarrow$ EM (Table)$\uparrow$ EM (Image)$\uparrow$ EM (Image & Table)$\uparrow$
@@
Table 10: Performance comparison of different model families fine-tuned on real and synthesized data on MultimodalQA. The ratios for EM scores and hallucination were calculated from filtered data (EM(Table) refers to the EM score calculated on samples that include the table modality). In the table, $\uparrow$ indicates that larger values are better (all EM values), while $\downarrow$ indicates that smaller values are better (hallucination rate). The scores for EM (Table) and EM (Image) may include samples that also contain other modalities. Larger models and those fine-tuned on synthesized data generally exhibit improved performance with reduced hallucination rates. In MultiModalQA, models demonstrate higher performance on questions involving the image modality compared to those involving tables.
@@
Table 11: Overall performance of different model families fine-tuned on real and synthesized data on WebQA. Hallucination is the proportion of hallucinated responses to incorrect answers. In this table, $\uparrow$ indicates that larger values are better, while $\downarrow$ indicates that smaller values are better. Fine-tuning on synthesized data consistently reduces hallucination rates and improves EM scores across all models, with larger models achieving the best performance. Unlike M2QA-Bench and MultModalQA, WebQA only includes image and text modality, as a result no EM(Image) and EM(Table) are reported.

### v5

@@ 5.3 Comparison with Human-Annotated Datasets
Unlike prior methods like MultiModalQA (Talmor et al., 2021) and WebQA (Chang et al., 2021), our approach is fully automated with no human involvement (minimal human evaluation was used in creating the M2QA-Bench only). This section compares the quality of our synthesized data against these human-annotated datasets. We trained LLaVA-1.6 (Liu et al., 2023b, a, 2024), InternVL-2 (Chen et al., 2023b, 2024c), and Idefics-2 (Laurençon et al., 2023, 2024b) on varying sizes of WebQA and MultiModalQA, evaluating on their respective test sets. We also trained the same models on FM2DS-generated training data and evaluated them on the same test sets to assess the effectiveness of the synthesized data in comparison to human-annotated datasets.
As shown in Table 4, models trained on FM2DS data outperform those trained on original datasets, despite using longer documents. WebQA seems easier than MultiModalQA, with better performance from fewer samples. On average, EM improved by 1.81 for MultiModalQA and 1.96 for WebQA using equal or fewer synthetic samples. While EM gains often lead to higher F1, some cases show F1 drops due to hallucinated answers reducing string overlap. Models trained on fewer synthetic samples from FM2DS match the performance of those trained on full datasets, showing faster convergence (see Section 5.4). Larger models also perform better with the same synthetic data; e.g., LLaVA-1.6-13B vs. 7B. GPT-4o leads among large LVLMs, likely due to its role in data generation (Refer to Appendix N for qualitative analysis).
When comparing these human annotated data with our synthesized data, a common question is: "Why not paraphrase existing datasets instead of synthesizing new ones?". The answer is paraphrasing does not enable domain adaptation, and as shown in Appendix L, fine-tuning on our synthetic data outperforms training on a paraphrased version of MultiModalQA.
@@ 5.4 Learning Efficiency Comparison
To evaluate learning efficiency, we ran experiments with InternVL-2-8B using incremental training sizes from 1k to 10k (in 1k steps) on both synthetic and human-annotated data. For real data, we used the same dataset for training and testing; e.g., when testing on WebQA, training samples were taken from the WebQA training set.
As shown in Figure 3(a & b), our synthesized data outperforms real data at smaller training sizes, though the gap narrows as sample size grows. Near 10k samples, learning efficiency with synthetic data declines more than with real data, likely due to its broader knowledge coverage. While this diversity aids early learning, it can lead to saturation, unlike real data, which offers more focused patterns and sustains steady learning (Hong et al., 2023; Maini et al., 2024).
In a related experiment on MultiModalQA, we used full Wikipedia pages via linked articles as training data instead of information snippets. This was not possible for WebQA, as its source links mostly point to WikiMedia pages with limited text. Figure 3(c) shows that models trained on full Wikipedia pages initially achieve better improvement per 1k samples; however, this advantage declines after approximately 3k samples. This suggests that, while full-page real data offers some early benefits over short snippets, it still lacks the generality and consistency of our synthesized dataset. In contrast, the synthesized data, with its built-in queries and diverse content, continues to support steady learning, likely due to its broader coverage and higher quality.
5.5 Cross-Dataset Evaluation
To assess the generalizability of our synthesized data, we conducted a cross-dataset evaluation using InternVL-2-8B. We used 1k samples from M2QA-Bench and 1k from the MultiModalQA test set, both with full Wikipedia pages as sources. The model was trained separately on 5k samples from our dataset and 5k from MultiModalQA, using the same input format. As shown in Table 5, the model trained on our data outperformed the one trained on MultiModalQA across both test sets, demonstrating stronger generalization. This also reflects the greater complexity and diversity of our benchmark, compared to MultiModalQA’s template-based questions. See Appendix J to see the performance on model trained on FM2DS’s data when using on a out of domain MMQA dataset.
@@ 5.6 Key Stages in FM2DS
Table 6: Results of the FM2DS with and without key steps like query generation and verification. All other steps are included in all of the results. The plus sign ("+") at the start each steps means all the previous steps were included as well.
To assess the impact of each data filtering and validation step, we evaluated InternVL-2-8B on 5k training samples under different configurations. As shown in Table 6, each step contributed to performance gains. Question validation improved relevance by filtering questions tailored to the complex, multimodal, multihop requirements of the test sets, and reduced hallucinations, boosting F1 with more accurate answers. Answer validation removed incorrect samples, while query generation distilled knowledge from larger models, improving EM and F1. Query validation reinforced consistency by ensuring proper structure. Refer to Appendix M for statistics like "Rejection Rate" on the validation stages used to remove specific samples during data synthesis.
@@ 5.7 M2QA-Bench Evaluation Results
To assess the complexity of M2QA-Bench and compare model performance, we evaluated a diverse set of models, as shown in Figure 4. GPT-4o stands out, outperforming even larger models like LLaMa-3.2-90B and Claude-3.5-Sonnet by a notable margin. Interestingly, smaller models in the 4B–8B range, particularly those from the Phi family (Abdin et al., 2024a, b), also achieve competitive results despite their scale. GPT-4o’s advantage may stem in part from its involvement in the initial data generation, but this remains an open question, as human annotators thoroughly reviewed and corrected the dataset to ensure fairness, reduce bias, and maintain high quality.
@@ 6 Human Evaluation of Answer Validation
To evaluate the accuracy and impact of our automatic answer validation component, we conducted a human evaluation study. After generating questions for 100 samples, we continued the pipeline using two methods: FM2DS with and without answer validation. These samples were divided into four batches of 25, with each batch evaluated by three participants to mitigate potential bias or human errors. Twelve participants used our evaluation platform (Appendix H) to judge answer correctness.
Participants were instructed to verify the correctness of each answer by reviewing the associated Wikipedia pages. Once they determined the accuracy of each answer, they were asked to select one of four options: (1) Method 1’s answer is correct, (2) Method 2’s answer is correct, (3) both methods generated the correct answer, or (4) neither answer is correct (Method 1 and Method 2 were randomly assigned to the datasets generated with and without answer validation, respectively.). The results, shown in Figure 5, reveal a clear trend: answer validation increases the likelihood of correct answers, even though the model without validation still produced correct answers in over 60% of cases. The Fleiss’ Kappa was 0.91, indicating strong agreement, though this is expected, as the task involved factual questions with definitive answers rather than subjective judgments.
@@ 7 Conclusion & Future Works
Limitations
While FM2DS offers a robust pipeline for synthesizing high-quality multimodal multihop QA data, several limitations remain.
First, although the framework incorporates strong validation steps, including factual consistency checks, named entity alignment, and hallucination detection, it is not immune to errors. Subtle factual inaccuracies and hallucinations may still persist, especially in answers grounded in complex visual content. Despite using multiple generations and automated agreement checks, there is still a risk that some incorrect samples pass through undetected.
Second, our reliance on large-scale generative models such as GPT-4o throughout multiple stages, including question generation, answer synthesis, captioning, and validation, makes the pipeline computationally expensive. This cost is further amplified by the need to regenerate failed samples that do not pass intermediate validation steps. In some settings, particularly when generating large-scale datasets, the repeated use of high-capacity models may pose practical limitations in terms of both time and resources.
Third, while our method improves factual accuracy and reduces hallucination, the validation pipeline is primarily designed for fact-based QA. This makes it less suitable for tasks involving subjective reasoning, commonsense inference, or open-ended discussion questions. Extending the pipeline to handle such cases would require fundamentally different validation strategies that go beyond factual grounding.
Finally, since the student models are trained on data generated by large models (used in both synthesis and supervision), there is a risk of knowledge leakage or model bias propagation. The synthetic data may overrepresent patterns and linguistic preferences from the teacher models, potentially limiting the generalizability of the student models trained on it.
@@
Potential Risks
The primary risk associated with this work lies in the possibility of propagating factual inaccuracies or biases through automatically synthesized data. While our validation pipeline aims to minimize hallucinations and ensure factual correctness, it may not catch all subtle errors. Additionally, overreliance on large language models for data generation could inadvertently reinforce biases encoded in those models. Our approach does not involve any sensitive personal data or downstream applications that could directly harm individuals.
Annotator Recruitment
To verify and refine the samples in our M2QA-Bench dataset, we recruited three human annotators with prior experience in NLP and data annotation (two men and one woman). These annotators were compensated fairly at a rate of \$25 per hour to reflect their expertise and time investment. All annotators were provided with detailed task descriptions and underwent an informed consent process prior to participation. The annotation process was conducted in accordance with ethical research guidelines and ensured voluntary participation and data confidentiality.
Evaluator Recruitment
For the human evaluation component of our study, we recruited twelve evaluators to compare answers generated with and without our validation pipeline. Evaluators were compensated at a rate of \$10 per hour and participated voluntarily after giving informed consent. They were clearly informed about the nature and purpose of the task. We ensured the task was low-risk, did not involve sensitive content, and that participation remained anonymous and non-intrusive.
Consent and Data Privacy
All participants in both annotation and evaluation roles were briefed on the nature of the task and explicitly consented to take part in the study. No personally identifiable information was collected or stored during any part of the research process. All data generated and reviewed by annotators and evaluators remained anonymous and was used strictly for academic research purposes.
Use of AI Assistants
We used AI assistants such as GitHub Copilot and ChatGPT to support coding, text editing, and formatting tasks during the development of this paper and the implementation of our framework. These tools were employed to accelerate workflow and refine writing, but all conceptual, experimental, and analytical decisions were made by the authors. We ensured that no sensitive data was provided to these tools during usage.
@@ Appendix A Prompts
Here are the documents: [Documents]
Here are examples: [Example(s)]
@@
Question: [Question]
Here are the documents: [Documents]
Here are examples: [Example(s)]
@@ A.3 Query Generation
Question: [Question]
Answer: [Answer]
Here are the documents: [Documents]
@@ Appendix B Experimental Settings
Appendix C Cost of Sample Generation
We analyze the cost of generating high-quality samples across different approaches. Using GPT-4o, the average cost of producing one high-quality sample is approximately \$0.035. By comparison, LLaMA 3.2-90B achieves similar quality at the cost of 42 H100 GPU hours for generating 5k samples (see Table 8 in Appnedix E), which is competitive with unimodal data synthesis methods. This aligns with prior work on unimodal data generation Chen et al. (2024b).
In contrast, human-written samples such as those in MMQA are substantially more expensive: each sample costs around \$2 and requires approximately 5 minutes of human effort. This comparison highlights the scalability and cost-effectiveness of large language models for multimodal data synthesis, offering orders-of-magnitude savings over manual annotation while maintaining comparable quality.
Appendix D The Effect of the Number of In-Context Examples
@@
Table 7 presents the results of evaluating the Intervl-2-8B model with varying numbers of in-context examples on the MultiModalQA and WebQA datasets. In the zero-shot setting, FM2DS exhibits limited understanding of multimodal multi-hop question answering, and occasionally circumvents the validation step by simply generating a question that is not multihop. For example, "Looking at the image of the Eiffel Tower, what engineering innovation allowed it to surpass previous structures in height?" prompts the model to use the image, but the answer is available in the page’s text on tall structures. As we move from zero-shot to one-shot, there is a significant boost in EM and F1 scores, reflecting improved performance with minimal context. The improvement from one-shot to two-shot is marginal, suggesting diminishing returns. With three in-context samples, the gains become minimal, indicating that additional samples beyond two provide little benefit. This diminishing return may stem from the model’s limited context window, which restricts its ability to fully utilize large in-context samples (Kaplan et al., 2020; Bertsch et al., 2024).
Appendix E Impact of Data Synthesis LLM Choice
@@
Model Setting MultimodalQA WebQA M2QA-Bench
LLaVA-1.6-13B FT -17.4 -14.2 -18.6
LLaVA-1.6-13B PT -22.7 -19.6 -24.1
InternVL-2-8B FT -19.1 -16.3 -20.2
InternVL-2-8B PT -25.3 -21.0 -26.8
GPT-4o PT -15.9 -13.1 -17.3
[…]
@@
InternVL-2.5-8B 5k/5k 72.13 76.82 80.46 85.91 5k/5k 79.27 84.74 86.39 91.18
InternVL-2.5-8B 10k/10k 75.41 78.92 83.73 88.16 10k/10k 82.34 86.95 89.57 93.23
InternVL-2.5-26B 5k/5k 76.98 81.27 85.15 90.32 5k/5k 85.62 89.79 91.37 94.56
InternVL-2.5-26B 10k/10k 80.42 83.84 88.29 92.17 10k/10k 88.48 91.53 93.68 95.83
InternVL-3-9B 5k/5k 74.72 79.45 82.63 88.03 5k/5k 83.19 87.57 89.42 93.14
InternVL-3-9B 10k/10k 77.36 81.71 85.18 89.94 10k/10k 86.27 89.62 92.18 95.28
InternVL-3-38B 5k/5k 82.64 86.28 90.73 94.34 5k/5k 89.94 91.93 94.27 96.35
InternVL-3-38B 10k/10k 85.18 88.31 92.86 95.73 10k/10k 91.16 93.83 95.62 97.14
@@
InternVL-2.5-8B None 63.27 70.16 None 69.38 77.42
InternVL-2.5-26B None 70.35 76.82 None 76.18 83.28
InternVL-3-9B None 65.47 71.92 None 72.58 80.27
InternVL-3-38B None 75.13 81.35 None 80.64 87.46
@@
Appendix F Role of Supporting Context
An important question in multimodal multihop QA is whether large language models can answer questions directly in a zero-shot setting without access to supporting context. To investigate this, we evaluated a zero-shot baseline where models were only prompted with the question and no additional multimodal context. As shown in Table 9, performance dropped significantly across models compared to settings where full context was provided.
Pre-trained models were the most affected by missing context. Fine-tuned models experienced smaller declines, indicating that FM2DS training equips them with reasoning strategies that generalize beyond explicit evidence. Notably, GPT-4o, despite not being fine-tuned on FM2DS, showed the smallest losses. This highlights GPT-4o’s strong inherent reasoning abilities, while also reinforcing that fine-tuning on FM2DS fosters similar robustness even when no supporting context is available.
Moreover, the drop is consistently larger on M2QA-Bench  highlighting that its tasks depend more heavily on fine-grained contextual grounding. This underscores both the higher complexity of M2QA-Bench and the central role of FM2DS in training models that remain resilient even when explicit multimodal evidence is absent.
Appendix G Investigating Performance of More Models on FM2DS Sythesized Data
Appendix H Human Evaluation Details
@@
In addition to these selections, participants had the option to provide a brief rationale for their choices. Although they have not been investigated for this research, these optional feedbacks were encouraged, as they offer valuable insights for qualitative analysis and potential future improvements in answer validation accuracy. The combination of structured and open-ended responses enhances the robustness of our evaluation and offers a more comprehensive view of user judgments, which we may explore in future iterations of our data synthesis methodology.
The evaluators had diverse academic and professional backgrounds, including graduate students in computer science, data science researchers, and software engineers with experience in NLP and machine learning. All evaluators were proficient in English and had prior familiarity with Wikipedia-style content and fact-based question answering tasks. This diversity contributed to reliable judgment across a wide range of topics and ensured that participants had the necessary background to assess factual correctness and relevance accurately. In total, twelve individuals participated in the evaluation: seven men and five women.
Appendix I Additional Statistics & Information on M2QA-Bench
@@
To further examine the diversity of questions in our benchmark—which also reflects the overall characteristics of the data generated by FM2DS—we conducted a 2D t-SNE analysis of question embeddings using ModernBERT (Warner et al., 2024). We sampled 500 questions each from M2QA-Bench, MMQA, WebQA, and ScienceQA (Lu et al., 2022b). ScienceQA serves as a fully human-authored dataset, while MMQA and WebQA primarily use templated questions. As shown in Figure 9, MMQA and WebQA display the least diversity. In contrast, M2QA-Bench, which includes questions generated from FM2DS, demonstrates greater similarity to human-generated data, reflecting a reduced domain gap and improved diversity compared to MMQA and WebQA.
Model PMC-VQA (Acc) SPIQA (F1) LegalBench (Acc)
GPT-4o (3-shot) 57.13 64.84 77.67
InternVL2-8B (FT on 5k) 65.72 77.91 83.45
Table 11: Cross-domain evaluation on health (PMC-VQA), science (SPIQA), and law (LegalBench). Fine-tuning InternVL-2-8B with 5k FM2DS samples substantially improves performance compared to GPT-4o (3-shot) and pretrained InternVL-2-8B baselines.
Appendix J FM2DS Generalizability to Other Domains
FM2DS can generate domain-specific synthetic data using just three in-context examples, enabling even small LVLMs to handle specialized multimodal multihop QA. As shown in Figure 7, FM2DS’s data (including M2QA-Bench) spans a wide range of domains. To further assess its generalizability, we trained InternVL-2-8B on 5k synthesized samples and evaluated it across three out-of-domain benchmarks: the health-related PMC-VQA (Zhang et al., 2023), the scientific benchmark SPIQA Pramanick et al. (2024), and the legal reasoning dataset LegalBench Guha et al. (2023).
Table 11 compares performance against GPT-4o (3-shot prompting) and the InternVL-2-8B fine-tuned on 5k samples generated by FM2DS for each specific domain. The fine-tuned model consistently outperforms GPT-4o across health, science, and law, showing clear improvements in every case. This demonstrates that FM2DS enables models to generalize effectively beyond the training distribution and can be readily adapted to build strong domain-expert systems across diverse fields.
Appendix K M2QA-Bench Examples
Figure 10: Multimodal and multihop reasoning example from M2QA-Bench where the model answers a question about the photograph "Raising the Flag on Iwo Jima" by synthesizing information from linked documents through a hyperlink, leveraging both visual and tabular data to determine the number of casualties from the USA.
Figure 11: Multimodal multihop reasoning example from M2QA-Bench where the model compares the release dates of two albums, "Music from Big Pink" and "Imagine," using textual and visual cues. The documents are connected through their shared topic, "music," and the answer is determined as the title of the earlier-released album.
@@
Appendix L Synthesizing Data vs. Paraphrasing Existing Human Annotated Datasets
Figure 12: Performance comparison of InternVL-2-8B trained on 1k samples from three settings: original MultimodalQA, paraphrased MultimodalQA (reworded using GPT-4o), and fully synthesized data from FM2DS. While paraphrasing existing questions yields only modest gains, our synthesized samples lead to significantly higher performance, highlighting the value of generating diverse and structurally novel multihop multimodal questions.
Paraphrasing questions from existing datasets introduces surface-level linguistic changes but preserves the original semantic intent and reasoning pathways, offering only marginal improvements in model training. In contrast, the data synthesized by FM2DS is intentionally crafted to introduce diverse question structures, span multiple domains, and require varied types of reasoning, pushing models toward more comprehensive multimodal understanding. To compare these approaches, we trained InternVL-2-8B using 1k samples from three settings: (I) the original MultimodalQA dataset, (II) a paraphrased version of MultimodalQA where questions were reworded using GPT-4o with the prompt "Please paraphrase the following question: [Question]" and (III) synthesized samples generated by FM2DS. For all conditions, we used full Wikipedia documents as sources. Figure 12 presents the results, showing that while paraphrasing provides a slight improvement, synthesizing new, high-quality samples with FM2DS leads to a substantial performance gain.
Stage Average Rejection Rate Average Rejection Rate
Question Validation 131 11.58%
Answer Validation 76 7.06%
Query Validation 58 5.48%
Appendix M Statistics on Usages of Each Validation Stage of FM2DS
As described in Section 3 and illustrated in Figure 2, FM2DS incorporates multiple validation stages to enhance data quality. It is essential to analyze how frequently each stage rejects the initially generated outputs. Table 12 presents statistics based on generating 1,000 examples using GPT-4o. Among the stages, question validation has the highest rejection rate, suggesting that this step is the most challenging. This may be because generating a question requires the model to synthesize all relevant knowledge and fully grasp the context. In contrast, answer validation benefits from the guidance provided by the question, making the task relatively easier. Query validation appears to be even more straightforward, as it primarily involves formatting the reasoning steps, something the model has effectively done during answer generation. Additionally, the use of question-specific image captions during answer generation likely contributes to a lower error rate by helping the model locate the correct information only text modality.
Appendix N Qualitative Analysis
@@
This analysis was conducted for 100 samples from each of the following benchmarks: (1) M2QA-Bench, (2) MultiModalQA, and (3) WebQA. The results are presented in Tables 13, 14, and 15. The responses generated by different models were analyzed across these datasets, focusing on the following metrics:
@@
Table 13: Performance of models fine-tuned on real vs. synthesized data on M2QA-Bench. EM scores and hallucination rates are computed on filtered data (hallucination = hallucinated responses / incorrect answers). $\uparrow$ indicates higher is better (EM), and $\downarrow$ indicates lower is better (hallucination). EM (Table) and EM (Image) may overlap with other modalities. Larger models and those trained on synthesized data achieve higher EM and lower hallucination, with table questions generally easier than image ones.
[…]
@@
Table 14: Performance of models fine-tuned on real vs. synthesized data on MultimodalQA. EM scores and hallucination rates are computed on filtered data (EM(Table) = EM on samples with table modality). $\uparrow$ means higher is better (EM), $\downarrow$ means lower is better (hallucination). EM (Table) and EM (Image) may overlap with other modalities. Larger models and those trained on synthesized data achieve higher EM with fewer hallucinations, with image-based questions generally easier than table-based ones.
@@
Table 15: Performance of models fine-tuned on real vs. synthesized data on WebQA. Hallucination is measured as hallucinated responses over incorrect answers. $\uparrow$ means higher is better (EM), $\downarrow$ means lower is better (hallucination). Fine-tuning on synthesized data improves EM and reduces hallucination across all models, with larger models performing best. Unlike M2QA-Bench and MultimodalQA, WebQA has only image and text, so EM(Image) and EM(Table) are not reported.